### v1

@@ Abstract
In a Jacobi–Davidson (JD) type method for singular value decomposition (SVD) problems, called JDSVD, a large symmetric and generally indefinite correction equation is approximately solved iteratively at each outer iteration, which constitutes the inner iterations and dominates the overall efficiency of JDSVD. In this paper, a convergence analysis is made on the minimal residual (MINRES) method for the correction equation. Motivated by the results obtained, a preconditioned correction equation is derived that extracts useful information from current searching subspaces to construct effective preconditioners for the correction equation and is proved to retain the same convergence of outer iterations of JDSVD. The resulting method is called inner preconditioned JDSVD (IPJDSVD) method. Convergence results show that MINRES for the preconditioned correction equation can converge much faster when there is a cluster of singular values closest to a given target, so that IPJDSVD is more efficient than JDSVD. A new thick-restart IPJDSVD algorithm with deflation and purgation is proposed that simultaneously accelerates the outer and inner convergence of the standard thick-restart JDSVD and computes several singular triplets of a large matrix. Numerical experiments justify the theory and illustrate the considerable superiority of IPJDSVD to JDSVD.
@@ 1 Introduction
(1.1) $A=U\Sigma V^{T},$
[…]
@@
Given a target $\tau>0$ that is not equal to any singular value of $A$, assume that the singular values of $A$ are labeled by
(1.2) $|\sigma_{1}-\tau|<\dots<|\sigma_{\ell}-\tau|<|\sigma_{\ell+1}-\tau|\leq\dots\leq|\sigma_{N}-\tau|.$
@@
We are interested in the $\ell$ singular triplets $(\sigma_{i},u_{i},v_{i})$, $i=1\dots,\ell$. Throughout the paper, we suppose that $\tau$ lies in the interval $(\sigma_{\min},\sigma_{\max})$ with $\sigma_{\min}$ and $\sigma_{\max}$ the smallest and largest singular values of $A$, respectively. Specifically, if $\tau$ is neither close to $\sigma_{\min}$ nor to $\sigma_{\max}$, then $(\sigma_{i},u_{i},v_{i})$, $i=1\dots,\ell$ are called interior singular triplets of $A$; otherwise they are called extreme, i.e., the largest or smallest, singular triplets. For some applications of extreme and interior singular triplets, see [17].
Projection methods [4, 21, 22, 26, 27] for large eigenvalue problems have been adapted to the computation of a partial SVD of $A$. There are four kinds of projection methods for computing approximate singular triplets of $A$: the standard extraction [2, 6, 7, 8, 14, 29], the harmonic extraction [7, 8, 9, 13, 18, 20], the refined extraction [8, 9, 10, 14, 28, 29], and the refined harmonic extraction [10, 15, 28, 29].
@@
For all projection methods on the computation of partial SVDs, a common core task is the construction and expansion of left and right searching subspaces so that they contain as much information on the desired singular vectors as possible. There are mainly three types of approaches to construct searching subspaces for the large SVD computation. Next we review two well-known approaches, and leave out the very new but complicated one that is used in the FEAST SVDsolvers [16, 17], which compute the singular triplets with all the singular values in a given interval. The first approach is the Lanczos bidiagonalization (LBD) process [2, 4]. It reduces $A$ to a sequence of small bidiagonal matrices and generates two sets of orthonormal bases of the left and right Krylov subspaces formed by $AA^{T}$ and $A^{T}A$ with relevant starting vectors. Such searching subspaces generally favor the left and right singular vectors associated with the extreme singular values of $A$; that is, the distances between these singular vectors and the corresponding Krylov subspaces generally tend to zero fast as the LBD process proceeds. As a consequence, combined with appropriate extraction approaches, state-of-the-art LBD type algorithms [8, 14, 15, 18, 19, 24] can compute the extreme singular triplets of $A$ efficiently. However, the LBD type methods do not suit well for the computation of interior singular values of $A$.
@@
The second approach is that used in Jacobi–Davidson (JD) type methods. The JD method proposed by Sleijpen and van der Vorst [23] for large eigenvalue problems is extended by Hochstenbach [7, 8] to the large SVD computation, referred to as JDSVD method. Wu and Stathopoulos [29] present a hybrid two-stage JDSVD method, called PHSVD. At the first stage, PHSVD transforms the SVD problem of $A$ into the mathematically equivalent eigenproblem of the cross-product matrix $A^{T}A$ and computes the desired singular values and right singular vectors with relatively low accuracy by some projection method. At the second stage, in order to obtain more accurate approximate singular triplets of $A$, it implements the JD method on the eigenproblem of the augmented $\begin{bmatrix}\begin{smallmatrix}\bm{0}&A\\ A^{T}&\bm{0}\end{smallmatrix}\end{bmatrix}$ by using the approximate singular vectors obtained in the first stage as an initial guess. Wu, Romero and Stathopoulos [28] develop the PRIMME__\__SVDS software that implements the PHSVD method with standard, refined and refined harmonic extractions.
@@
At each step of JDSVD and the second stage of PHSVD, one must solve some large shifted linear system, called correction equation, that involves $\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ and the current approximate singular triplet, and uses its solution to expand the searching subspaces. The correction equations are approximately solved iteratively, called the inner iterations, while the outer iterations compute approximate singular triplets with respect to the searching subspaces. The least solution accuracy requirement on the correction equations and convergence speed of the inner iterations critically affect the outer convergence and dominate the overall efficiency of JDSVD, respectively. The least solution accuracy means that the outer iterations of the resulting inexact JDSVD methods behave as if all the correction equations were solved exactly. The authors [10] have figured out such accuracy requirement and proved that it generally suffices to solve all the correction equations with low or modest accuracy, i.e., the relative errors $\varepsilon\in[10^{-4},10^{-2}]$ of approximate solutions.
@@
As will turn out, the correction equations involved in JDSVD methods are generally indefinite and may be ill conditioned, which is particularly true when the desired singular values are clustered at $\tau$; the correction equation is positive definite only when $\sigma_{1}=\sigma_{\max}$ is of interest. As a result, it is generally very hard to iteratively solve them even with low accuracy. To this end, for the overall efficiency of JDSVD, effective preconditioning correction equations is absolutely necessary. For a general large linear system $Bx=z$, the preconditioning is to transform it into a mathematically equivalent one that is easier to be solved by a suitable iterative solver, e.g., the conjugate gradient (CG) method, the minimal residual (MINRES) method, the generalized MINRES (GMRES) method, and the biconjugate gradient stabilized (BiCGstab) method [1, 5, 22]. There are a lot of preconditioning procedures [22], and among others two major classes of popular algebraic preconditioning techniques are incomplete LU factorization (ILU)-based preconditioners $LU\approx B$ and sparse approximate inverse (SAI) preconditioners $C\approx B^{-1}$.
@@
In JDSVD-type methods, since the coefficient matrices in the correction equations involve $\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ and are generally symmetric indefinite, the ILU and SAI precondioning procedures encounter severe difficulties and are not appropriate choices in several aspects. For instance, (i) constructing an SAI preconditioner is costly. In a JDSVD method with adaptively changing shifts, the coefficient matrix in the correction equation changes at each outer iteration, which requires that one construct a new SAI preconditioner from scratch. Therefore, it may be unaffordable to do so. (ii) ILU preconditioners are generally ineffective and have intrinsic numerical instability for indefinite symmetric or highly nonsymmetric systems. (iii) The symmetry of correction equations enables us to fully exploit the merits of MINRES, but ILU and SAI preconditioners destroy the symmetry, so that we have to use a much inferior nonsymmetric solver GMRES or BiCGstab to solve the preconditioned unsymmetric correction equations. Numerically, we have observed that BiCGstab with ILU preconditioners is very often inefficient and performs even worse than the unpreconditioned MINRES, in terms of total inner iterations and total CPU time [10]. Therefore, in the context of JDSVD, other effective preconditioning approaches for the correction equation are extremely appealing.
@@
For symmetric positive definite and indefinite linear systems, the CG and MINRES are preferences among Krylov iterative solvers, respectively. Setting aside initial vectors, their convergence is uniquely determined by the eigenvalues of an underlying coefficient matrix. These two methods converge fast when the eigenvalues are clustered, and they may converge very slowly when the coefficient matrix has relatively small eigenvalues in magnitude [5]. For such a linear system, favorable preconditioning is to preserve the symmetry of the preconditioned linear system and, meanwhile, shrink the eigenvalue inclusion interval of the preconditioned definite matrix or the two intervals containing negative and positive eigenvalues of the preconditioned indefinite coefficient matrix. For the correction equation in JDSVD, as will be clear, an ideal preconditioning procedure is to annihilate the $m$ clustered smallest eigenvalues in magnitude of the coefficient matrix that are well separated from the other eigenvalues, so that the eigenvalue distribution of the preconditioned coefficient matrix is favorable and MINRES converges fast. Our careful investigation will show that this is equivalent to forming a new, called preconditioned, correction equation, in which the orthogonal projector is now formed by the left and right singular vectors $u_{i}$ and $v_{i}$ associated with the $m$ clustered singular values $\sigma_{i},i=1,2,\ldots,m$ rather than only $\sigma_{1}$. Stacking these vectors, we obtain an $m$-dimensional eigenspace of the coefficient matrix in the correction equation associated with its $m$ clustered smallest eigenvalues in magnitude, which will be shown to eliminate the effects of these smallest eigenvalues so that MINRES converges much faster than it does for the original one.
@@
We have observed from extensive numerical experiments that, as the outer iterations of JDSVD proceed, the current subspaces also contain increasingly richer information on the left and right singular vectors $u_{i}$ and $v_{i}$, $i=1,\ldots,m$ corresponding to the $m$ clustered singular values. These reasonably good approximate singular vectors form a good approximate eigenspace of the coefficient matrix in the correction equation corresponding to its $m$ clustered smallest eigenvalues in magnitude. We will fully exploit such rich information to derive a symmetric preconditioned correction equation that eliminates the effects of those smallest eigenvalues in magnitude of the coefficient matrix, so that MINRES can be accelerated greatly.
@@
We first make a detailed convergence analysis on MINRES for the correction equation in JDSVD for computing one singular triplet $(\sigma_{1},u_{1},v_{1})$. Then we consider the case that there are $m$ clustered singular values closest to the target $\tau$, and show that MINRES may converge very slowly. For the preconditioned correction equation, we prove an extremely important property: its solution has the same quality as the solution to the standard correction equation when expanding the current subspaces; that is, the solution to the preconditioned correction equation expands the subspaces as effectively as the solution to the standard one and, therefore, and does not spoil the subsequent outer iterations. Then we establish convergence results on MINRES for the preconditioned correction equation, and prove that it can converge substantially faster than it does for the standard one. In computations, we shall set up some necessary criteria and propose an adaptive approach to pick up appropriate approximate singular triplets to form the preconditioned correction equation. The resulting method is called the inner-preconditioned JDSVD (IPJDSVD) method. On the other hand, importantly and practically, we propose a new thick-restart IPJDSVD algorithm with deflation and purgation that computes $\ell$ singular triplets of $A$. Whenever $m>k_{\min}$, the initial subspace dimension of the standard thick-restart, the new thick-restart is more effective than the standard one and saves outer iterations.
@@
Since the goal of this paper is to propose an effective preconditioning procedure for the correction equation in JDSVD and a new thick-restart strategy, we shall only consider the standard extraction-based JDSVD method and the shifts in all the correction equations are $\tau$. However, the theoretical results and analysis on the correction equations and preconditioning are directly applicable to the harmonic, refined and refined harmonic extraction-based JDSVD methods and their adaptively-changing-shift variants [10], so are the corresponding thick-restart algorithms.
In section 2 we review the standard extraction-based JDSVD method. In section 3 we make a convergence analysis on MINRES for the correction equation in JDSVD. In section 4 we derive the preconditioned correction equation, and extend the results and analysis in section 3 to the preconditioned correction equation. We consider how to dynamically form a preconditioned correction equation in computations. In section 5 we propose the thick-restart IPJDSVD with deflation and purgation to compute $\ell$ singular triplets of $A$. Numerical experiments are reported in section 6 to illustrate the high performance of our preconditioning procedure and the thick-restart IPJDSVD. We conclude the paper in section 7.

### v2

@@ Abstract
In a Jacobi–Davidson (JD) type method for singular value decomposition (SVD) problems, called JDSVD, a large symmetric and generally indefinite correction equation is solved iteratively at each outer iteration, which constitutes the inner iterations and dominates the overall efficiency of JDSVD. In this paper, a convergence analysis is made on the minimal residual (MINRES) method for the correction equation. Motivated by the results obtained, at each outer iteration a new correction equation is derived that extracts useful information from current subspaces to construct effective preconditioners for the correction equation and is proven to retain the same convergence of outer iterations of JDSVD. The resulting method is called inner preconditioned JDSVD (IPJDSVD) method; it is also a new JDSVD method, and any viable preconditioner for the correction equations in JDSVD is straightforwardly applicable to those in IPJDSVD. Convergence results show that MINRES for the new correction equation can converge much faster when there is a cluster of singular values closest to a given target. A new thick-restart IPJDSVD algorithm with deflation and purgation is proposed that simultaneously accelerates the outer and inner convergence of the standard thick-restart JDSVD and computes several singular triplets. Numerical experiments justify the theory and illustrate the considerable superiority of IPJDSVD to JDSVD, and demonstrate that a similar two-stage IPJDSVD algorithm substantially outperforms the most advanced PRIMME_SVDS software nowadays for computing the smallest singular triplets.
@@ 1 Introduction
(1) $A=U\Sigma V^{T},$
where the superscript $T$ denotes the transpose of a matrix, $U=[u_{1},\dots,u_{M}]\in\mathbb{R}^{M\times M}$ and $V=[v_{1},\dots,v_{N}]\in\mathbb{R}^{N\times N}$ are orthogonal, and $\Sigma=\operatorname{diag}\{\sigma_{1},\dots,\sigma_{N}\}\in\mathbb{R}^{M\times N}$ with $\sigma_{i}\geq 0,i=1,\dots,N$. The $(\sigma_{i},u_{i},v_{i})$, $i=1,\dots,N$ are the singular triplets of $A$ with $\sigma_{i}$ the singular values and $u_{i},v_{i}$ the associated left and right singular vectors.
@@
Given a target $\tau\geq 0$ that is not equal to any singular value of $A$, assume that the singular values of $A$ are labeled by
(2) $|\sigma_{1}-\tau|<\dots<|\sigma_{\ell}-\tau|<|\sigma_{\ell+1}-\tau|\leq\dots\leq|\sigma_{N}-\tau|.$
@@
We are interested in the $\ell$ singular triplets $(\sigma_{i},u_{i},v_{i})$, $i=1\dots,\ell$. Specifically, if $\tau$ is inside $(\sigma_{\min},\sigma_{\max})$ but neither close to $\sigma_{\min}$ nor to $\sigma_{\max}$, then $(\sigma_{i},u_{i},v_{i})$, $i=1\dots,\ell$ are called interior singular triplets of $A$; otherwise they are called extreme, i.e., the largest or smallest, singular triplets. Particularly, $\tau=0$ corresponds to the smallest singular triplets and $\tau$ closer to $\sigma_{\max}$ than to the second largest singular value corresponds to the largest singular triplets.
@@
For the concerning large SVD problem, only projection methods can be used. Over the years, projection methods [3, 20, 21, 25, 26] for large eigenvalue problems have been adapted to the computation of a partial SVD of $A$. Given left and right searching subspaces, there are four classes of projection methods for computing approximate singular triplets of $A$, which are based on the standard extraction [1, 5, 6, 7, 13, 28], the harmonic extraction [6, 7, 8, 12, 17, 19], the refined extraction [7, 8, 9, 13, 27, 28], and the refined harmonic extraction [9, 14, 27, 28], respectively. For these methods, a common core task is the construction of left and right searching subspaces so that they contain as much information on the desired singular vectors as possible. Besides a very recent approach that is used in the FEAST SVDsolvers [15, 16] for computing the singular triplets with all the singular values in a given interval, there are mainly two types of approaches for this. The first approach is the Lanczos bidiagonalization (LBD) process [1, 3]. It reduces $A$ to a sequence of small bidiagonal matrices and generates two sets of orthonormal bases of the left and right Krylov subspaces formed by $AA^{T}$ and $A^{T}A$ with relevant starting vectors. Such searching subspaces generally favor the extreme left and right singular vectors, and the distances between these singular vectors and the corresponding Krylov subspaces generally tend to zero fast as the LBD process proceeds. Combined with appropriate extraction approaches, state-of-the-art LBD type algorithms [13, 14, 17, 18, 22] can compute the extreme singular triplets of $A$ efficiently if the desired singular values are not many and are well separated. Unfortunately, the smallest singular values from applications are often clustered, so that the LBD type methods may work poorly.
@@
The second approach is that used in Jacobi–Davidson (JD) type methods proposed by Hochstenbach [6] for the large SVD computation, referred to as JDSVD method. Wu and Stathopoulos [28] present a hybrid two-stage JDSVD method, called PHSVD. At the first stage, PHSVD transforms the SVD problem of $A$ into the mathematically equivalent eigenproblem of the cross-product matrix $A^{T}A$ and computes the desired singular values and right singular vectors to best achievable accuracy by a chosen projection method efficiently. Then if further accuracy is required, it switches to the second stage that solves the eigenproblem of the augmented $\begin{bmatrix}\begin{smallmatrix}\bm{0}&A\\ A^{T}&\bm{0}\end{smallmatrix}\end{bmatrix}$ by a chosen projection method using the approximate singular vectors obtained in the first stage as initial guesses. Wu, Romero and Stathopoulos [27] develop the PRIMME__\__SVDS software programmed in C language; it implements the PHSVD method with the refined extraction in the second stage and uses JDQMR by default, while the method in the first stage is taken as the default ‘DYNAMIC’ that dynamically switches between JDQMR and the GD+k method to minimize the time, where QMR in JDQMR means that a symmetric QMR (sQMR) is used to solve the correction equations [23]. A great merit of the hybrid PRIMME__\__SVDS is that, regarding the outer iterations, JDQMR or GD+k in the first stage is generally much faster than it is in the second stage and, in the meantime, sQMR is generally faster than it is in the second stage, which is true when the smallest singular values are desired. An advantage of sQMR over the MINRES method [4, 21] is that sQMR allows a symmetric indefinite preconditioner and uses the same matrix-vector products at each iteration as MINRES; without preconditioner, sQMR is mathematically equivalent to MINRES [23].
@@
In the sequel, we always assume that PRIMME__\__SVDS uses JDQMR in the second stage. Then at each step of JDSVD [6] and the second stage of PRIMME__\__SVDS, one must solve a certain correction equation that involves $\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ and the current approximate singular triplet, and uses its solution to expand the searching subspaces. Whenever the current approximate singular value starts to converge with fair accuracy, it will replace $\tau$ in the correction equation so that its solution provides richer information on the desired left and right singular vectors. Such kind of correction equations is approximately solved iteratively, called the inner iterations, while the outer iterations compute approximate singular triplets with respect to the searching subspaces. The least solution accuracy requirement on the correction equations and convergence speed of the inner iterations critically affect the outer convergence and dominate the overall efficiency of JDSVD, respectively. The least solution accuracy means that the outer iterations of the resulting inexact JDSVD methods behave as if all the correction equations were solved exactly. The authors in [9] have figured out such accuracy requirement and proven that it generally suffices to solve all the correction equations with the relative errors $\varepsilon\in[10^{-4},10^{-2}]$ of approximate solutions.
@@
As will be seen, the correction equations may be highly indefinite and/or ill conditioned, which is definitely true when the desired singular values are the smallest ones and/or clustered. As a result, it is generally very hard to iteratively solve them using Krylov solvers. To this end, for the overall efficiency of a JDSVD-type method, effectively preconditioning correction equations plays a crucial role. For a general large linear system $Bx=z$, the preconditioning is to transform it into a mathematically equivalent one that is easier to be solved by a suitable Krylov solver, e.g., the conjugate gradient (CG) method, the minimal residual (MINRES) method, the generalized MINRES (GMRES) method, and the biconjugate gradient stabilized (BiCGstab) method [4, 21]. Among others, two major classes of popular algebraic preconditioning techniques are incomplete LU factorization (ILU)-based preconditioners $LU\approx B$ and sparse approximate inverse (SAI) preconditioners $C\approx B^{-1}$ [21].
@@
Since the coefficient matrices in the correction equations involve $\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ and are generally symmetric indefinite, the ILU and SAI preconditioning procedures encounter severe difficulties and are generally not appropriate choices in several aspects. For instance, (i) constructing an SAI preconditioner is costly. In a JDSVD-type method with adaptively changing shifts, the coefficient matrix in the correction equation changes at each outer iteration, which requires that one construct a new SAI preconditioner from scratch. (ii) ILU preconditioners are generally ineffective and have intrinsic numerical instability for indefinite symmetric or highly nonsymmetric systems. (iii) The symmetry of correction equations enables us to fully exploit the merits of MINRES, but ILU and SAI preconditioners destroy the symmetry, so that we have to use a much inferior nonsymmetric solver GMRES or BiCGstab to solve the preconditioned unsymmetric correction equations. Although sQMR allows an indefinite preconditioner and may have convergence behavior similar to MINRES, it indeed has irregular convergence and exhibits plateaus [23, 28], impairing the overall efficiency of JDQMR.
@@
For symmetric positive (or negative) definite and indefinite linear systems, the CG and MINRES are preferences among Krylov iterative solvers, respectively. Setting aside initial vectors, their convergence is uniquely determined by the eigenvalues of an underlying coefficient matrix. The methods converge fast when the eigenvalues are clustered, and they may converge very slowly when the coefficient matrix has relatively small eigenvalues in magnitude [4]. For such a linear system, we seek to preserve the symmetry of the preconditioned linear system and, meanwhile, shrink the eigenvalue inclusion interval(s) of the preconditioned coefficient matrix. Starting with this perspective, for the correction equation in JDSVD, we will see that an ideal preconditioning procedure is to annihilate the $m$ clustered smallest eigenvalues in magnitude of the coefficient matrix that are well separated from the other eigenvalues, so that the eigenvalue distribution of the preconditioned coefficient matrix is favorable and MINRES converges fast. This will turn out to amount to deriving a preconditioned correction equation and forming a new correction equation at each outer iteration, in which the orthogonal projector is formed by the left and right singular vectors $u_{i}$ and $v_{i}$ associated with the $m$ clustered singular values $\sigma_{i},i=1,2,\ldots,m$ rather than only $\sigma_{1}$. Stacking these vectors, we obtain an $m$-dimensional eigenspace of the coefficient matrix in the correction equation associated with its $m$ clustered smallest eigenvalues in magnitude, which will eliminate the effects of these eigenvalues on MINRES and thus accelerate MINRES substantially.
@@
We have observed from extensive numerical experiments that, as the outer iterations of JDSVD proceed, the current subspaces provide increasingly better approximations to the left and right singular vectors $u_{i}$ and $v_{i}$, $i=1,\ldots,m$ corresponding to the $m$ clustered singular values. These approximations form a good approximate eigenspace of the coefficient matrix corresponding to its $m$ clustered smallest eigenvalues in magnitude. We will fully exploit such rich information to derive a new symmetric correction equation, which is equivalent to preconditioning the original one in JDSVD and eliminates the effects of these smallest eigenvalues as much as possible, so that MINRES can converge much faster. The resulting method is called the inner-preconditioned JDSVD (IPJDSVD) method. Importantly, from another perspective, IPJDSVD can be considered to be a new JDSVD method with a new correction equation solved at each outer iteration, so that any applicable preconditioner to the correction equations in JDSVD can be straightforwardly used for the new correction equations in IPJDSVD.
@@
In this paper, we first make a detailed convergence analysis on MINRES for the correction equation in JDSVD for computing one singular triplet $(\sigma_{1},u_{1},v_{1})$. Then we consider the case that there are $m$ clustered singular values closest to the target $\tau$, and show that MINRES may converge very slowly. For the preconditioned correction equation, we prove an extremely important property: its solution has the same quality as the solution to the standard correction equation when expanding the current subspaces; that is, the solution to the preconditioned correction equation expands the subspaces as effectively as the solution to the standard one and, therefore, does not spoil the subsequent outer iterations. Then we establish convergence results on MINRES for the preconditioned correction equation, and prove that it can converge substantially faster than it does for the standard one. In computations, we shall set up some necessary criteria and propose an adaptive approach to pick up appropriate approximate singular triplets to form the preconditioned correction equation. Importantly and practically, we propose a new thick-restart IPJDSVD algorithm with deflation and purgation that computes $\ell$ singular triplets of $A$. Whenever $m>k_{\min}$, the initial subspace dimension of the standard thick-restart, the new thick-restart is more effective than the standard one and saves outer iterations.
@@
Since the goal of this paper is to propose an effective preconditioning procedure for the correction equation in JDSVD and a new thick-restart strategy, we shall take the standard extraction-based JDSVD method as example and the shifts in all the correction equations as the fixed $\tau$. However, the theoretical results and analysis on the correction equations and preconditioning are adaptable to the harmonic, refined and refined harmonic extraction-based JDSVD methods and their adaptively-changing-shift variants [9], so are the corresponding thick-restart algorithms.
In Section 2 we review the standard extraction-based JDSVD method. In Section 3 we make a convergence analysis on MINRES for the correction equation in JDSVD. In Section 4 we derive the preconditioned correction equation, and extend the results and analysis in Section 3 to the preconditioned correction equation. We consider how to dynamically form a preconditioned correction equation in computations. In Section 5 we propose the thick-restart IPJDSVD with deflation and purgation to compute $\ell$ singular triplets of $A$. Numerical experiments are reported in Section 6 to illustrate the high performance of the thick-restart IPJDSVD. In the meantime, to compare our algorithm with the currently most advanced PRIMME__\__SVDS software [27] in C language, in the same spirit we have also developed a hybrid two-stage thick-restarted IPJDSVD algorithm, called IPJDSVD_HYBRID: in the first stage, a similar IPJDSVD method computes the desired largest or smallest singular triplets based on the eigenproblem of $A^{T}A$ (or $AA^{T}$ if $m<n$) with the same outer stopping tolerance as that in the first stage of PRIMME_SVDS, and it then switches to the second stage whenever necessary, in which IPJDSVD is based on the eigenproblem of $\begin{bmatrix}\begin{smallmatrix}\mathbf{0}&A\\ A^{T}&\mathbf{0}\end{smallmatrix}\end{bmatrix}$ using the approximate singular triplets computed at the first stage as initial guesses, as done and detailed in this paper, and further improves approximate singular triplets until it converges for a reasonably prescribed stopping tolerance. We demonstrate that IPJDSVD_HYBRID substantially outperforms PRIMME_SVDS when computing the simple smallest singular triplets and is at least competitive with the latter when computing the simple largest ones. Section 7 concludes the paper.